# State Entropy Regularization for Robust Reinforcement Learning

### v2

@@ 1 Introduction
Recent works islam2019marginalized ; seo2021state ; yuan2022renyi ; kim2023accelerating ; bolland2024off proposed to regularize the standard RL objective with the entropy of the distribution over states induced by a policy Hazan2018ProvablyEM , either alone or in combination with policy entropy. This has empirically shown better exploration and consequently, improved sample efficiency (seo2021state, ). However, to our knowledge, formal studies on the effect of state entropy regularization and its connection with robustness do not exist in the literature yet. Thus, an open question is whether robustness is a by-product of state entropy regularization just like it is for policy entropy, and if it is, what type of robustness results from state entropy regularization?
@@ 2 Preliminaries
=${\mathcal{R}}=\{\tilde{r}\not>r\mid D(r,\tilde{r})\leq\epsilon\},$ (1)
where $\epsilon>0$, $D$ is interpreted as a dissimilarity measure between reward functions and $\tilde{r}\not>r$ means that there exists some $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$ such that $\tilde{r}(s,a)\leq r(s,a)$ (to ensure the adversary can degrade performance rather than only improve it). Another perspective on robust RL comes from considering the existence of an adversary who is perturbing the nominal MDP to minimize the performance of the policy $\pi$. Thus, we refer to $\epsilon$ as the adversary’s budget for perturbations.
@@
where $\alpha>0$ is a hyperparameter controlling the regularization strength. In recent works (islam2019marginalized, ; seo2021state, ; yuan2022renyi, ; kim2023accelerating, ; bolland2024off, ), regularization with the entropy of the state distribution has been considered, either standalone or in combination with policy entropy, yielding
@@
Risk-averse RL. While robust RL tackles the epistemic uncertainty in RL, in which the uncertainty lies on the parameters of the MDP, risk-averse RL garcia2015comprehensive aims at the aleatoric uncertainty caused by the inherent randomness of the MDP. In this context, we seek robustness to unfavorable realizations of the return $R(\tau):=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$ where $\tau=(s_{0},a_{0},r_{0},\ldots)$ is sampled from $p^{\pi}_{P}(\tau):=\mathbb{P}[s_{0},a_{0},r_{0},\ldots|\pi,P]$. While several risk-averse formulations exist, a common objective rockafellar2000optimization ; chow2015risk is given by the Conditional Value-at-Risk (CVaR) of the return, formally defined as:
where $\beta\in(0,1)$ is a percentile and $(x)^{+}:=\max(x,0)$.
@@ 3.1 Reward-Robustness
where the reward uncertainty set is ~$\tilde{\mathcal{R}}^{\pi}(\epsilon,\alpha):=\left\{\tilde{r}\mid\textsc{lse}_{{\mathcal{S}}}(\frac{r^{\pi}-\tilde{r}^{\pi}}{\alpha})\leq\frac{\epsilon}{\alpha}+\log(S)\right\}$. Moreover, the worst-case reward chosen by the adversary is given by $\tilde{r}(s,a)=r(s,a)-\frac{1}{A}\frac{\log(d^{\pi}_{P}(s))}{\pi(a\mid s)}-\epsilon.$
@@
Figure 2: Uncertainty sets ~$\tilde{\mathcal{R}}^{\pi}_{\epsilon}(\alpha)$ induced by state entropy regularization in a two-state MDP with $r^{\pi}(s_{1})=1$, $r^{\pi}(s_{2})=0$, and $\epsilon=\frac{1}{2}$. The $x$- and $y$-axes correspond to the perturbed rewards $\tilde{r}^{\pi}(s_{1})$ and $\tilde{r}^{\pi}(s_{2})$, respectively. As $\alpha$ increases, the uncertainty set expands, interpolating between $\ell_{\infty}$- and $\ell_{1}$-type constraints in the limits $\alpha\to 0$ and $\alpha\to\infty$.
We compare the obtained uncertainty sets and robustness guarantees with policy entropy and state-action entropy regularization in Table 1. The comparison extends previous formulations that neglect either the temperature parameter $\alpha$ or the adversary’s budget $\epsilon$ eysenbach2022maximum ; brekelmans2022your . More insight into the proofs of these derivations is included in Appx. A.2. As shown in the adversarial reward column, policy entropy regularization depends on the policy $\pi(\cdot|s)$ only. Thus, it tackles adversaries with access to local information of each state. In contrast, state entropy regularization depends on the state occupancy distribution $d^{\pi}_{P}$, enabling robustness against adversaries that exploit global properties of the agent’s behavior across the state space. Consequently, policies regularized via state entropy are inherently more resistant to global perturbations, while policy entropy provides protection primarily against local, state-specific variations. This difference analytically confirms our initial intuition.
@@
~$\tilde{{\mathcal{R}}}^{\pi}_{\epsilon}(\alpha)=\left\{\tilde{r}\mid\mathrm{LSE}_{{\mathcal{S}}}\left(\left(\frac{r^{\pi}-\tilde{r}^{\pi}}{\alpha}\right)^{\alpha}\right)-\alpha\log(S)\leq\epsilon\right\}.$
Then, the mapping $\alpha\mapsto\tilde{{\mathcal{R}}}^{\pi}_{\epsilon}(\alpha)$ is non-increasing in the sense that $\alpha_{2}\geq\alpha_{1}$ implies ~$\tilde{\mathcal{R}}(\epsilon,\alpha_{1})\subseteq\tilde{\mathcal{R}}(\epsilon,\alpha_{2})$. Furthermore, the limit sets satisfy
@@ 3.2 Lower Bound under Kernel Uncertainty
$\tilde{{\mathcal{P}}}^{\pi}(\epsilon,\alpha):=\bigg{\{}\tilde{P}\mid\alpha\log\bigg{(}\frac{1}{{\mathcal{S}}}\sum_{s\in{\mathcal{S}}}\bigg{(}\frac{d^{\pi}_{P}(s)}{d^{\pi}_{\tilde{{\mathcal{P}}}}(s)}\bigg{)}^{\frac{1}{\alpha}}\bigg{)}\leq\epsilon\bigg{\}}$
@@ 4.1 Entropy Regularization Cannot Fully Solve Kernel Robustness
We consider a general class of kernel robust RL problems in which the uncertainty set is defined as a function of the nominal transition kernel $P$ and the policy $\pi$, denoted $\tilde{\mathcal{P}}(P,\pi)$. This captures a wide range of uncertainty models, including $\ell_{p}$-constrained and rectangular uncertainty sets commonly studied in robust RL l_pKernelRobust ; derman2021twice . We now state a more general claim: Any regularization function that depends only on the nominal kernel and the policy cannot, in general, characterize the robust return over such uncertainty sets. The following impossibility result formalizes this observation.
@@
Let $\Omega(\pi,P)\not\equiv 0$ be a regularization function. There does not exist an uncertainty set $\tilde{\mathcal{P}}(P,\pi)$ such that $\min_{\tilde{P}\in\tilde{\mathcal{P}}}\mathbb{E}_{\rho^{\pi}_{\tilde{P}}}[r]=\mathbb{E}_{\rho^{\pi}_{P}}[r]+\Omega(\pi,P).$
@@ 4.2 Entropy Regularization and Risk-Averse Performance
We denote the optimal policy for the unregularized objective by $\pi^{*}$ and the optimal policy for the entropy-regularized objective (whether state or policy entropy) by $\pi^{*}_{R}$.
@@
For every temperature $\alpha>0$, constant $M\in\mathbb{R}$, and $10<\beta<$ 1, there exists an MDP $\mathcal{M}$ such that:
@@ 4.3 On the Sample Sensitivity of State Entropy Regularization
For the statistical explanation, we assume that the state entropy is estimated using a $k$-nearest neighbor estimator (singh2003nearest, ) in a continuous state space $\mathbb{R}^{D}$, as it is common in the related literature mutti2021task ; liu2021behavior ; seo2021state . For fixed $k$, under mild regularity conditions on the MDP—such as the smoothness of $d^{\pi}_{P}$ for all the policies $\pi$—the estimation error of the state entropy decays at a rate $O(n^{-1/D})$, [^2] For simplicity, here we consider the i.i.d. case in which the entropy estimate is computed with a single data point coming from each trajectory, sampled with a geometric distribution with parameter $\gamma$. where $n$ denotes the number of sampled trajectories (SinghKNN, ). In contrast, both the policy entropy and expected return can be estimated with error decaying at the standard rate $O(n^{-1/2})$ pmlr-v206-metelli23a . As a result, in high-dimensional settings, state entropy estimation introduces a dominant source of statistical error, contributing to the increased sample sensitivity of entropy-regularized methods. In practice, we note that the effective dimensionality $D$ is often reduced by computing entropy over a subset of state features (as in Section 5) or by estimating entropy over latent state representations, as demonstrated by seo2021state .
@@ 5 Experiments
We evaluate the robustness properties of state entropy regularization using a set of representative environments and perturbation types. These are designed to test three core hypotheses: (1) State entropy regularization offers protection against spatially correlated severe perturbations to the dynamics; (2) State entropy regularization induces robustness to $\ell_{1}$ reward uncertainty; (3) The robustification of state entropy regularization is sensitive to the number of rollouts used during training. We validate these claims in both discrete (MiniGrid chevalier2018minimalistic ) and continuous (Mujoco todorov2012mujoco ) domains. We compare policies trained without regularization, with policy entropy regularization, and with state entropy regularization. Note that the state-entropy regularized policies are trained with some policy entropy regularization to favor training stability ahmed2019understanding . [^3] A discussion of the interplay between state entropy and policy entropy regularization is in Appx. B.4. For all the types, we select the largest regularization coefficients that degrade nominal performance less than 5%. To ensure a fair comparison, the methods are trained with the same base algorithm—A2C mnih2016asynchronous for MiniGrid and PPO schulman2017proximal for Mujoco.
@@
Transition kernel robustness via entropy regularization.  To evaluate how well each regularization method handles spatially localized perturbations to the transition dynamics, we construct environments where specific regions of the state space are blocked off or obstructed. These changes are considered “catastrophic” because they remove critical paths or create bottlenecks that fundamentally alter the agent’s ability to reach the goal using its original policy. In MiniGrid, we randomly place a large wall segment that blocks a connected area of the maze (see Figure 4(b) for the perturbed environment). This constitutes a major local disturbance, as it affects transitions between nearby states and introduces a sharp structural change in the environment. As shown in Figure 4(b), the agent trained with state entropy regularization significantly outperforms both policy entropy and unregularized baselines. We note that we report results for all horizon lengths to avoid misleading conclusions based on a single fixed task length. In cases where one method clearly dominates across the full horizon range, we can draw decisive conclusions. When the curves intersect or remain close, it indicates that the relative performance is more sensitive to task structure. To verify that this effect extends to continuous control, we conduct a similar experiment in the Pusher environment, where we place a wall between the puck and the goal. Again, as shown in Figure 5(a), state entropy regularization yields substantially more robust performance.
@@
Reward robustness.   To evaluate how different forms of entropy regularization handle reward uncertainty, we consider a MiniGrid variant in which the reward location is randomized at test time. This introduces structured variation in the reward function—formally resembling an $\ell_{1}$-bounded uncertainty—where the total change in reward mass is limited but can be concentrated arbitrarily across the state space. The key intuition is that if the reward happens to lie along the agent’s typical trajectory, then policy entropy regularization—by injecting stochasticity at the action level—may suffice to recover it. However, when the reward is placed in less frequently visited areas, broader state coverage becomes necessary. In such cases, state entropy regularization offers a natural advantage by encouraging the agent to visit a wider range of states. As shown in Figure 4(d), policy entropy improves performance relative to the unregularized baseline, but adding state entropy yields a more substantial gain, particularly when the reward is far from the nominal goal location.
@@
Figure 5: Pusher experiments and ablations. As in the discrete experiment we train all agents on a nominal task then introduce a local catastrophe by blocking off the optimal path, visualized in Fig. 1. All plots report success rate defined as when the pucks final distance from the goal is smaller than 0.1. We evaluate each policy for 200 episodes and show 96% confidence bounds across 25 seeds.

### v3

@@ 1 Introduction
Recent works islam2019marginalized; seo2021state; yuan2022renyi; kim2023accelerating; bolland2024off proposed to regularize the standard RL objective with the entropy of the distribution over states induced by a policy Hazan2018ProvablyEM, either alone or in combination with policy entropy. This has empirically shown better exploration and consequently, improved sample efficiency (seo2021state). However, to our knowledge, formal studies on the effect of state entropy regularization and its connection with robustness do not exist in the literature yet. Thus, an open question is whether robustness is a by-product of state entropy regularization just like it is for policy entropy, and if it is, what type of robustness results from state entropy regularization?
@@ 2 Preliminaries
We note that, while the above formulation assumes finite MDPs, the main results extend to continuous state or action spaces under mild regularity assumptions. In that case, sums over states and actions are replaced by integrals with respect to the corresponding densities, and the discrete entropy is replaced by the differential entropy of the occupancy measure.
@@
${\mathcal{R}}=\{\tilde{r}\not>r\mid D(r,\tilde{r})\leq\epsilon\},$ (1)
where $\epsilon>0$, $D$ is interpreted as a dissimilarity measure between reward functions defining the notion of proximity within the uncertainty set and $\tilde{r}\not>r$ means that there exists some $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$ such that $\tilde{r}(s,a)\leq r(s,a)$ (to ensure the adversary can degrade performance rather than only improve it). Another perspective on robust RL is to interpret the minimization in ${\mathcal{J}}(\pi,{\mathcal{P}},{\mathcal{R}})$ as the action of an adversary that perturbs the nominal MDP within the uncertainty set so as to minimize the policy’s expected return. Accordingly, $\epsilon$ quantifies the adversary’s budget for allowable perturbations.
@@
where $\alpha>0$ is a hyperparameter controlling the regularization strength. In recent works (islam2019marginalized; seo2021state; yuan2022renyi; kim2023accelerating; bolland2024off), regularization with the entropy of the state distribution has been considered, either standalone or in combination with policy entropy, yielding
@@ 3.1 Reward-Robustness
where the reward uncertainty set is $\tilde{\mathcal{R}}^{\pi}(\epsilon,\alpha):=\left\{\tilde{r}\mid\textsc{lse}_{{\mathcal{S}}}(\frac{r^{\pi}-\tilde{r}^{\pi}}{\alpha})\leq\frac{\epsilon}{\alpha}+\log(S)\right\}$. Moreover, the worst-case reward chosen by the adversary is given by $\tilde{r}(s,a)=r(s,a)-\frac{1}{A}\frac{\log(d^{\pi}_{P}(s))}{\pi(a\mid s)}-\epsilon.$
@@
Figure 2: Uncertainty sets $\tilde{\mathcal{R}}^{\pi}_{\epsilon}(\alpha)$ induced by state entropy regularization in a two-state MDP with $r^{\pi}(s_{1})=1$, $r^{\pi}(s_{2})=0$, and $\epsilon=\frac{1}{2}$. The $x$- and $y$-axes correspond to the perturbed rewards $\tilde{r}^{\pi}(s_{1})$ and $\tilde{r}^{\pi}(s_{2})$, respectively. As $\alpha$ increases, the uncertainty set expands, interpolating between $\ell_{\infty}$- and $\ell_{1}$-type constraints in the limits $\alpha\to 0$ and $\alpha\to\infty$.
We compare the obtained uncertainty sets and robustness guarantees with policy entropy and state-action entropy regularization in Table 1. The comparison extends previous formulations that neglect either the temperature parameter $\alpha$ or the adversary’s budget $\epsilon$ eysenbach2022maximum; brekelmans2022your. More insight into the proofs of these derivations is included in Appx. A.2. As shown in the adversarial reward column, policy entropy regularization depends on the policy $\pi(\cdot|s)$ only. Thus, it tackles adversaries with access to local information of each state. In contrast, state entropy regularization depends on the state occupancy distribution $d^{\pi}_{P}$, enabling robustness against adversaries that exploit global properties of the agent’s behavior across the state space. Consequently, policies regularized via state entropy are inherently more resistant to spatially-structured perturbations, while policy entropy provides protection primarily against local, state-specific variations. This difference analytically confirms our initial intuition.
@@
$\tilde{{\mathcal{R}}}^{\pi}_{\epsilon}(\alpha)=\left\{\tilde{r}\mid\mathrm{LSE}_{{\mathcal{S}}}\left(\left(\frac{r^{\pi}-\tilde{r}^{\pi}}{\alpha}\right)^{\alpha}\right)-\alpha\log(S)\leq\epsilon\right\}.$
Then, the mapping $\alpha\mapsto\tilde{{\mathcal{R}}}^{\pi}_{\epsilon}(\alpha)$ is non-increasing in the sense that $\alpha_{2}\geq\alpha_{1}$ implies $\tilde{\mathcal{R}}(\epsilon,\alpha_{1})\subseteq\tilde{\mathcal{R}}(\epsilon,\alpha_{2})$. Furthermore, the limit sets satisfy
@@ 3.2 Lower Bound under Kernel Uncertainty
$\tilde{{\mathcal{P}}}^{\pi}(\epsilon,\alpha):=\bigg\{\tilde{P}\mid\alpha\log\bigg(\frac{1}{{\mathcal{S}}}\sum_{s\in{\mathcal{S}}}\bigg(\frac{d^{\pi}_{P}(s)}{d^{\pi}_{\tilde{P}}(s)}\bigg)^{\frac{1}{\alpha}}\bigg)\leq\epsilon\bigg\}$
@@ 4.1 Entropy Regularization Cannot Fully Solve Kernel Robustness
A kernel-robust RL problem is defined by a function $\tilde{\mathcal{P}}(\mathrm{MDP},\pi)$ which maps the MDP and policy parameters to an uncertainty set—that is, the set of possible transition kernels from which the adversary may choose. We consider a class of kernel robust RL problems in which the uncertainty set function depends only on the nominal transition kernel $P$ and the policy $\pi$. This captures a wide range of uncertainty models, including $\ell_{p}$-constrained and rectangular uncertainty sets commonly studied in robust RL l_pKernelRobust; derman2021twice. We now state a more general claim: Any regularization function that depends only on the nominal kernel and the policy cannot, in general, characterize the robust return over such uncertainty set functions. The following impossibility result formalizes this observation.
@@
Let $\Omega(\pi,P)\not\equiv 0$ be a regularization function. There does not exist an uncertainty set function $\tilde{\mathcal{P}}:(P,\pi)$ such that for every MDP holds:
$\min_{\tilde{P}\in\tilde{\mathcal{P}}}\mathbb{E}_{\rho^{\pi}_{\tilde{P}}}[r]=\mathbb{E}_{\rho^{\pi}_{P}}[r]+\Omega(\pi,P).$ (7)
@@ 4.2 Entropy Regularization and Risk-Averse Performance
While robust RL primarily addresses epistemic uncertainty—uncertainty about the parameters of the MDP—risk-averse RL  garcia2015comprehensive instead targets aleatoric uncertainty, which arises from the inherent randomness of the MDP itself. To quantify risk-averse performance, we use the common Conditional Value at Risk (CVaR) objective  rockafellar2000optimization; chow2015risk, denoted $\mathrm{CVaR}_{\beta}$, which measures the expected return in the worst $\beta$-fraction of outcomes. We denote the optimal policy for the unregularized objective by $\pi^{*}$ and the optimal policy for the entropy-regularized objective (whether state or policy entropy) by $\pi^{*}_{R}$.
@@
For every temperature $\alpha>0$, constant $M\in\mathbb{R}$, and $0<\beta<1$, there exists an MDP $\mathcal{M}$ such that:
@@ 4.3 On the Sample Sensitivity of State Entropy Regularization
For the statistical explanation, we assume that the state entropy is estimated using a $k$-nearest neighbor estimator (singh2003nearest) in a continuous state space $\mathbb{R}^{D}$, as it is common in the related literature mutti2021task; liu2021behavior; seo2021state. For fixed $k$, under mild regularity conditions on the MDP—such as the smoothness of $d^{\pi}_{P}$ for all the policies $\pi$—the estimation error of the state entropy decays at a rate $O(n^{-1/D})$, [^2] For simplicity, here we consider the i.i.d. case in which the entropy estimate is computed with a single data point coming from each trajectory, sampled with a geometric distribution with parameter $\gamma$. where $n$ denotes the number of sampled trajectories (SinghKNN). In contrast, both the policy entropy and expected return can be estimated with error decaying at the standard rate $O(n^{-1/2})$ pmlr-v206-metelli23a. As a result, in high-dimensional settings, state entropy estimation introduces a dominant source of statistical error, contributing to the increased sample sensitivity of entropy-regularized methods. In practice, we note that the effective dimensionality $D$ is often reduced by computing entropy over a subset of state features (as in Section 5) or by estimating entropy over latent state representations, as demonstrated by seo2021state.
@@ 5 Experiments
We evaluate the robustness properties of state entropy regularization using a set of representative environments and perturbation types. These are designed to test three core hypotheses: (1) State entropy regularization offers protection against spatially correlated severe perturbations to the dynamics; (2) State entropy regularization induces robustness to $\ell_{1}$ reward uncertainty; (3) The robustification of state entropy regularization is sensitive to the number of rollouts used during training. We validate these claims in both discrete (MiniGrid chevalier2018minimalistic) and continuous (Mujoco todorov2012mujoco) domains. We compare policies trained without regularization, with policy entropy regularization, and with state entropy regularization. Note that the state-entropy regularized policies are trained with some policy entropy regularization to favor training stability ahmed2019understanding. [^3] A discussion of the interplay between state entropy and policy entropy regularization is in Appx. B.5. For all the types, we select the largest regularization coefficients that degrade nominal performance less than 5%. To ensure a fair comparison, the methods are trained with the same base algorithm—A2C mnih2016asynchronous for MiniGrid and PPO schulman2017proximal for Mujoco.
@@
Transition kernel robustness via entropy regularization.  To evaluate how well each regularization method handles spatially correlated perturbations to the transition dynamics, we construct environments where specific regions of the state space are blocked off or obstructed. These changes are considered “catastrophic” because they remove critical paths or create bottlenecks that fundamentally alter the agent’s ability to reach the goal using its original policy. In MiniGrid, we randomly place a large wall segment that blocks a connected area of the maze (see Figure 4(b) for the perturbed environment). This constitutes a major spatially-structured disturbance, as it affects transitions between nearby states and introduces a sharp structural change in the environment. As shown in Figure 4(b), the agent trained with state entropy regularization significantly outperforms both policy entropy and unregularized baselines. We note that we report results for all horizon lengths to avoid misleading conclusions based on a single fixed task length. In cases where one method clearly dominates across the full horizon range, we can draw decisive conclusions. When the curves intersect or remain close, it indicates that the relative performance is more sensitive to task structure. To verify that this effect extends to continuous control, we run analogous experiments in Mujoco’s Pusher and Ant environments by placing obstacles along the optimal trajectories, fully blocking them. Again, as shown in Figs 5(a), 6(c) state entropy regularization yields substantially more robust performance as its induced behavioral diversity enables bypassing these local catastrophes.
To highlight the challenge of learning policies that are robust to spatially localized perturbations, we compare against the PRMDP and NRMDP frameworks Chen2019, two popular robustness baselines in continuous control. Both methods formulate robustness as a two-player zero-sum game: the agent selects actions while an adversary perturbs the actions. As shown in Figure  5(a), the agent trained with state entropy regularization significantly outperforms this baseline. While we do not expect state entropy to consistently outperform robust RL methods across all types of perturbations, this result—along with our analysis in the main paper—suggests that it is particularly effective against spatially correlated perturbations, which are not the primary focus of most standard approaches.
@@
Reward robustness.   To assess how different entropy regularization methods handle reward uncertainty, we randomize the reward location at evaluation time in both MiniGrid and Pusher. In MiniGrid, a new goal location is sampled uniformly from the entire grid. In Pusher, the goal is relocated to a random position at distance $r$ from its nominal location. This setup induces structured variation in the reward function, formally resembling an $\ell_{1}$-bounded uncertainty—where the total change in reward mass is limited but can be concentrated arbitrarily across the state space. The key intuition is that if the reward happens to lie along the agent’s typical trajectory, then policy entropy regularization—by injecting stochasticity at the action level—may suffice to recover it. However, when the reward is placed in less frequently visited areas, broader state coverage becomes necessary. In such cases, state entropy regularization offers a natural advantage by encouraging the agent to visit a wider range of states. As shown in Figures 4(d), 5 (b) policy entropy improves performance relative to the unregularized baseline, but adding state entropy yields a more substantial gain, particularly when the reward is far from the nominal goal location.
@@
Figure 5: Pusher experiments and ablations. Like in the discrete experiment, we train all agents on a nominal task, then introduce a localized catastrophe by blocking off the optimal path, visualized in Fig. 1. All plots report success rate defined as when the pucks final distance from the goal is smaller than 0.1. We evaluate each policy for 200 episodes and show 96% confidence bounds across 25 seeds.
@@
Figure 6: Ant experiment. (a) Visualization of the vertical wall perturbation. (b) All methods perform similarly on the nominal task. (c) Success rate in the perturbed environment, i.e., the fraction of evaluation runs in which the ant successfully crosses the obstacle. We evaluate each policy for 200 episodes and show 96% confidence bounds across 25 seeds.